OpenStudy (s117):

Anita has 180 minutes to prepare mailers for her job. It takes her at least 15 minutes to set up materials and 3 minutes per mailer to prepare them. If m represents the number of mailers she prepares, then 3m + 15 ≤ 180 represents the number she can finish. What is the greatest number of mailers Anita can prepare in the time she has?

OpenStudy (anonymous):

Solve the equation. \(3m + 15 \le 180\) Subtract 15 to both sides, what's 180 - 15?

OpenStudy (anonymous):

*inequality

OpenStudy (s117):



OpenStudy (s117):

165

OpenStudy (anonymous):

Yes, now we have: \(3m \le 165\) Now divide 3 to both sides.

OpenStudy (anonymous):

165 / 3 = ?

OpenStudy (s117):

55

OpenStudy (anonymous):

Yes, so we have: \(m \le 55\) So the greatest number of mailers Anita can have is 55.

OpenStudy (s117):

thnx @iGreen.